# Direct $N$-body simulations of rotating and extremely massive Population III star clusters

### 31 Mar 2026 v1

@@ Abstract
Conclusions. We confirm two of the hypothesized formation channels of galactic nuclei seed black holes: gravitational runaway mergers of black holes and of Population III stars, which core-collapse into IMBHs thereafter. Higher initial star cluster bulk rotation correlates with earlier core collapse and, in the event counts reported here, with more coalescences/collisions and lower retained (compact) binary abundances. Initial bulk rotation is a primary control parameter of cluster evolution: faster rotation accelerates early angular-momentum transport, gravothermal collapse, mass segregation, and amplifies post-collapse expansion, which also favours the formation of a compact central IMBH subsystem.
@@ 1 Introduction
There are three main formation channels of seed black holes (BHs) for galactic nuclei that are typically considered today (see the reviews by Rees, 1984; Greene et al., 2020, and sources therein). (i) Gravitational runaway mergers between stars and compact objects happen throughout cosmic time in dense star clusters. They can be separated into a “fast” and a “slow” regime following Greene et al. (2020). In the fast regime (a couple of Myr from star cluster formation and natal gas expulsion), gravitational runaway mergers can happen during early star cluster evolution, when stars evolve and merge either through binary stellar evolution or dynamical collisions (Portegies Zwart & McMillan, 2002; Sakurai et al., 2017; Gieles et al., 2018; Reinoso et al., 2018, 2022; Wang et al., 2022; Vergara et al., 2023, 2025a; Rantala et al., 2025). The slow regime (around 100 Myr to Gyr from star cluster formation and natal gas expulsion) is populated by gravitational runaway mergers that occur between BHs. Gravitational runaway mergers of both kinds are postulated to produce intermediate-mass BHs (IMBHs) with masses of order $10^{2}-10^{4}$ $\mathrm{M}_{\odot}$. Both the fast and the slow regimes have been confirmed extensively by simulations of dense star clusters using various methods and both mechanisms are instrumental to growing IMBHs (e.g. Giersz et al., 2015; Rodriguez et al., 2019; Arca Sedda et al., 2019, 2020b, 2024; Di Carlo et al., 2020a, b, 2021; Rizzuto et al., 2021, 2022; Leveque et al., 2022a; Maliszewski et al., 2022; Vergara et al., 2025b; Paiella et al., 2025). (ii) Population III (Pop III) stars above the pair-instability mass gap can collapse directly into black hole seeds with masses of $\sim 10^{2}$ $\mathrm{M}_{\odot}$ (e.g. Bromm & Larson, 2004; Bromm, 2013; Woosley, 2017; Haemmerlé et al., 2020; Mestichelli et al., 2024). Within the Pop III cluster, stars can collide and form an even more massive star, thus collapsing into a heavier black hole seed (e.g. Katz et al., 2015; Sakurai et al., 2017; Reinoso et al., 2018, 2022; Vergara et al., 2021; Tanikawa et al., 2022a; Wang et al., 2022).Reinoso et al. (2025)shows that IMBHs with masses up to $10^{4}$ $\mathrm{M}_{\odot}$ are possible to form through stellar mergers and gas accretion. (iii) A pristine gas cloud can avoid fragmentation, thus collapsing directly into a massive star that eventually forms a heavy seed with masses of around $10^{4}$-$10^{6}$ $\mathrm{M}_{\odot}$ (e.g. Bromm & Loeb, 2003; Begelman et al., 2006, 2008; Begelman, 2010; Chon & Omukai, 2025; van Dokkum et al., 2025). Extremely massive stars ($M\sim 10^{5}-10^{6}\penalty 10000\ \mathrm{M}_{\odot}{}$) can explode by the general relativistic instability supernova mechanism (e.g. Shibata & Shapiro, 2002; Sakurai et al., 2015; Uchida et al., 2017).
@@
As noted above, metal-poor Pop III star clusters at high redshifts ($z\gtrsim 10$) with a top-heavy initial mass function (IMF) (e.g. Sharda & Krumholz, 2022) are prime candidates to produce black-hole seeds for galactic nuclei and nuclear star clusters (NSCs Neumayer et al., 2020; Greene et al., 2020; Askar et al., 2021, 2022; Schleicher et al., 2022). The IMBHs that form in these clusters can grow even more massive through tidal disruption events over long time-scales (Stone et al., 2017; Sakurai et al., 2019; Wang et al., 2025b). The single- and multi-generation mergers from the gravitational runaway process of BHs and other compact objects originating from Pop III star clusters will be relevant gravitational wave (GW) detection events (e.g. Schneider et al., 2000, 2002, 2003; Kinugawa et al., 2014, 2016, 2021a, 2021b; Hartwig et al., 2016; Belczynski et al., 2017; Ng et al., 2022; Mestichelli et al., 2024; Liu et al., 2024a, b; Wu et al., 2025; Wang et al., 2025a). Liu et al. (2025) shows that binary BH mergers in Pop III star clusters can produce GW, similar to those observed in GW231123. Plunkett et al. (2025) develops observational strategies to constrain the stellar demographics of Pop III stars with next-generation telescopes, especially for the proposed third-generation ground-based GW detectors, the Cosmic Explorer (CE Reitze et al., 2019; Evans et al., 2021), Einstein Telescope (ET Punturo et al., 2010b, a; Sathyaprakash et al., 2012; Maggiore et al., 2020; Liu & Bromm, 2020, 2021), and for future space-borne detectors operating in the mHz and deci-Hz frequency bands, like the laser interferometer space antenna (LISA Amaro-Seoane et al., 2013, 2017, 2023), (TianQin Luo et al., 2016; Liu et al., 2020), or the deci-Hz GW observatory (DECIGO Kawamura et al., 2011, 2021), where IMBHs are expected to be bright GW sources (see e.g. Miller & Hamilton, 2002; Amaro-Seoane et al., 2007; Amaro-Seoane, 2018; Jani et al., 2020; Arca Sedda et al., 2020a, 2021).
@@
Using detailed binary population synthesis (BPS), Tanikawa et al. (2021b) conducted studies on the merger rate density of Pop III binary BHs below, within, and above the pair-instability mass gap from isolated binary stellar evolution. They find that mergers between two low mass BHs (in their models, low mass means $M<50$ $\mathrm{M}_{\odot}$) independent of mass ratio and semi-major axis distributions of the primordial Pop III binaries could be identified from the observed BHBH mergers by (Advanced) Laser Interferometer Gravitational-Wave Observatory ((a)LIGO Aasi et al., 2015; Abbott et al., 2018, 2019), (Advanced) Virgo Interferometer ((a)Virgo Acernese et al., 2015; Abbott et al., 2018, 2019) and by extension also Kamioka Gravitational Wave Detector (KAGRA Abbott et al., 2018, 2020; Kagra Collaboration et al., 2019), although the predicted present-day (10 Gyr) merger rates would be comparatively low ($\sim 0.1\penalty 10000\ \mathrm{yr}^{-1}\mathrm{Gpc}^{-3}$) with regard to the merger rate density inferred by GWTC-4 catalogue of $15.4^{+6.0}_{-4.5}$ $\mathrm{yr}^{-1}\mathrm{Gpc}^{-3}$ (The LIGO Scientific Collaboration et al., 2025). Similarly, mergers between a low mass BH and a high mass BH (in their models, high mass means $M>130$ $\mathrm{M}_{\odot}$ due to the pair instability mass gap) or mergers between two high mass BHs will be detectable using the afore-mentioned currently available GW detectors according to Tanikawa et al. (2021b). However, the authors caution that if two conditions hold simultaneously, namely a wide minimum semi-major axis in primordial Pop III binaries due stellar expansion in the protostellar phases and fast stellar rotation that causes excitation of (quasi-)chemically homogeneous evolution of these stars, during which Pop III stars stay more compact than their non-rotating counterparts due to mixed Helium throughout the star (e.g. Maeder, 1987; Yoon & Langer, 2005; de Mink et al., 2009). Rapidly rotating Pop III stars (Nandal et al., 2024; Tsiatsiou et al., 2024) have been proposed as a possible source of the high nitrogen abundance observed in galaxies at high redshift (e.g., Bouwens et al., 2010; Tacchella et al., 2023; Marques-Chaves et al., 2024; Naidu et al., 2025). Rotation in Pop III stars induces strong internal mixing, which enlarges the convective core and results in brighter stars, thus improving the chances of detecting these primordial objects in the early universe with the JWST (Hassan et al., 2025).
@@
Direct detections of Pop III stars, their remnants, or their host clusters remain lacking because they lie at very high redshift. Schauer et al. (2022) reported a candidate extremely massive Pop III star at $z=6.2$, but this requires confirmation. Vanzella et al. (2020) claimed a Pop III complex at $z=6.629$ with the MUSE Deep Lensed Field targeting the Hubble Frontier Field (HFF) galaxy cluster MACS J0416. Using JWST NIRCam and lensing, Fujimoto et al. (2025) identified a strong Pop III cluster candidate at $z=6.5$. Another candidate is the high redshift galaxy RX J2129–z8He, which exhibits He II emission at $1640\mathring$$\mathrm${A}} (Wang et al., 2024). Despite JWST, unambiguous direct observations remain difficult (Rydberg et al., 2013; Katz et al., 2022). de Souza et al. (2013) suggest that several hundred JWST supernova detections could constrain the Pop III IMF. Future wide-field surveys with Euclid (Laureijs et al., 2011; Tanikawa et al., 2022b) and Roman will likely outperform JWST in constraining direct-collapse black holes above the pair-instability gap because they cover larger areas (Lazar & Bromm, 2022; Vikaeus et al., 2022).
@@
Because direct observations of Pop III stars and their environments are lacking, the parameter space for their formation is poorly constrained (e.g. Klessen, 2019; Klessen & Glover, 2023). Fraser et al. (2017) inferred a Pop III IMF from 29 extremely metal-poor Pop II stars which are expected to form in environments enriched by Pop III SNe, and found a Pop I/II-like IMF, but the small sample prevents strong conclusions. Theory and simulations generally predict top-heavy Pop III IMFs; such IMFs favour XRB populations with brighter and more spatially irregular X-ray emission (Sartorio et al., 2023). Hydrodynamic models are therefore widely used to probe plausible Pop III IMFs and their environmental dependence (e.g. Hirano & Bromm, 2017, 2018; Susa, 2019; Chon et al., 2021; Sharda et al., 2021; Latif et al., 2022). The binary fraction and initial binary statistics remain highly uncertain (Stacy & Bromm, 2013; Liu et al., 2021a) and, being environment-dependent, affect massive-star and compact-object demographics (Hirano et al., 2018; Sugimura et al., 2020; Kroupa, 2025).
@@
The observational and theoretical uncertainties are inherited by the initial conditions of Pop III star cluster simulations. One such parameter concerns the degree of initial Pop III star cluster rotation. A set of direct N-body simulations of initially rotating Pop III star clusters were performed by Vergara et al. (2021) ($N=10^{3}-10^{4}$ and evolution only up to $2\penalty 10000\ \mathrm{Myr}$). They used Miyamoto-Nagai models with flattening and rotation and found that not only the collision rate increases with increasing bulk rotation, but additionally the number of escapers is reduced the larger the initial rotation is. The low particle numbers, the short simulation time and other simplifying assumptions, such as neglecting stellar evolution make it difficult to generalize these results. In general, for simulations in collisional dynamics of rotating star clusters, distribution functions from, e.g., Goodman (1983); Longaretti & Lagoute (1996); Einsel & Spurzem (1999); Varri & Bertin (2012) are typically used. Kamlah et al. (2022b) recently ran simulations of Pop II star clusters with rotating King models from Einsel & Spurzem (1999) (using the 2-D Fokker-Planck code Fopax, Einsel & Spurzem, 1999; Kim et al., 2002, 2004, 2008). Apart from the formation and dissolution of a rotating bar of BHs that is related to the gravothermal-gravogyro catastrophe (Hachisu, 1979, 1982; Akiyama & Sugimoto, 1989; Einsel & Spurzem, 1999; Hong et al., 2013; Kamlah et al., 2022b), they found a possible dependence of BHBH binary abundances on initial star cluster bulk rotation; for a fast rotating model (rotating King model $W_{0}=6.0,\penalty 10000\ \omega_{0}=1.2$), they found significantly many more BHBH binaries during a phase of star cluster evolution when BHBH mergers in the aforementioned ”slow” regime are relevant (Greene et al., 2020). Similarly, Webb et al. (2019) found that increased initial star cluster rotation precipitates the formation of circularized BHBH binaries. Since Pop III are postulated to harbour many more (IM)BHs than Pop I star clusters, relative to their size, we test whether earlier findings apply to Pop III clusters.
@@
Furthermore, the lack of concrete observations of Pop III stars adds uncertainty to their stellar evolution. The most impactful parameter that differentiates stars of same masses in Pop I and Pop III populations is the metallicity (therefore, Pop III stars are also referred to as extreme metal poor (EMP) stars, e.g., in Tanikawa et al., 2020). The lower the metallicity, the weaker are radiation driven winds that affect massive stars. Generally, this statement also holds for pulsation-driven winds (Nakauchi et al., 2020). In the regime of $Z/\mathrm{Z}_{\odot}{}<10^{-4}$ both the winds become negligibly small. However, it has been suggested that Pop III stars form with very high rotation rates, when the magnetic fields are negligible and thus there is negligible magnetic braking (Stacy et al., 2011, 2013; Hirano & Bromm, 2018). The winds due to the weakened stellar magnetic field of such stars may be very powerful (Liu et al., 2021b; Jeena et al., 2023).
@@
The lower the metal in a Pop III star, the more drastic its internal evolutionary changes, and the lower its opacity (Ekström et al., 2008). First, the lower the metallicity of a star, the more compact it is, because the line-driven radiation pressure diminishes. Furthermore, most massive Pop I/II stars are characterized by a red supergiant (RSG) evolutionary phase with convective envelopes, while most Pop III stars end with a blue supergiant (BSG) phase that have radiative envelopes (Tanikawa et al., 2020). These properties affect binary stellar evolution. While Pop I/Pop II stars tend to undergo unstable mass transfer and common envelope evolution (CEE) in the RSG phase, BSGs from Pop III stars undergo stable mass transfer, so that less mass is ejected from the binary system. As a result, BHBH binaries from Pop III stellar populations can be more massive than Pop I/II counterparts even ignoring wind mass loss making Pop III star clusters a very attractive target of GW event detection and progenitor studies (e.g. Inayoshi et al., 2017; Kinugawa et al., 2021a, b; Tanikawa et al., 2021b, a, 2022c, 2022a; Liu et al., 2024a). Dense stellar environments substantially increase the fraction of mergers involving BHs with primary masses exceeding the pair-instability mass limit (Mestichelli et al., 2024, 2025). Furthermore, shorter initial orbital periods lead to higher merger rates Santoliquido et al. (2023), which can boost the production of GW sources in the early universe.
@@
In this paper, we present eight direct N-body simulations of extremely massive, rotating Pop III clusters, run with and without self-consistent general relativistic (GR) merger recoil kicks. The main novelty is the inclusion of initial bulk rotation. We examine how initial bulk rotation, state-of-the-art Pop III stellar-evolution models, GR merger recoil kicks (GRK), primordial binaries, a continuous IMF, and tidal-field mass loss jointly shape cluster dynamics and IMBH formation.
@@
The paper is structured as follows. Sect. 2 summarizes the methodology, the implementation of the GRK and the Pop III stellar evolution fitting formulae. Sect. 3 outlines the initial conditions. Sect. 4 presents the results. In Sect. 5 we summarize and discuss the work.

### 13 Apr 2026 v2

@@ Abstract
Conclusions. We confirm two of the hypothesized formation channels of galactic nucleus seed black holes: gravitational runaway mergers of black holes and of Population III stars, which core-collapse into IMBHs thereafter. A higher initial star cluster bulk rotation correlates with earlier core collapse and, in the event counts reported here, with more coalescences and collisions, as well as lower retained (compact) binary abundances. Initial bulk rotation is a primary control parameter of cluster evolution: faster rotation accelerates early angular-momentum transport, gravothermal collapse, mass segregation, and amplifies post-collapse expansion, which also favours the formation of a compact central IMBH subsystem.
@@ 1 Introduction
There are three main formation channels of seed black holes (BHs) for galactic nuclei that are typically considered today (see the reviews by Rees, 1984; Greene et al., 2020, and sources therein). (i) Gravitational runaway mergers between stars and compact objects happen throughout cosmic time in dense star clusters. They can be separated into a ‘fast’ and a ‘slow’ regime following Greene et al. (2020). In the fast regime (a couple of million years from star cluster formation and natal gas expulsion), gravitational runaway mergers can happen during early star cluster evolution, when stars evolve and merge either through binary stellar evolution or dynamical collisions (Portegies Zwart & McMillan, 2002; Sakurai et al., 2017; Gieles et al., 2018; Reinoso et al., 2018, 2022; Wang et al., 2022; Vergara et al., 2023, 2026; Rantala et al., 2025). The slow regime (around 100 Myr to a gigayear from star cluster formation and natal gas expulsion) is populated by gravitational runaway mergers that occur between BHs. Gravitational runaway mergers of both kinds are postulated to produce intermediate-mass BHs (IMBHs) with masses of order $10^{2}-10^{4}$ $\mathrm{M}_{\odot}$. Both the fast and the slow regimes have been confirmed extensively by simulations of dense star clusters using various methods and both mechanisms are instrumental to growing IMBHs (e.g. Giersz et al., 2015; Rodriguez et al., 2019; Arca Sedda et al., 2019, 2020b, 2024; Di Carlo et al., 2020a, b, 2021; Rizzuto et al., 2021, 2022; Leveque et al., 2022a; Maliszewski et al., 2022; Vergara et al., 2025; Paiella et al., 2025). (ii) Population III (Pop III) stars above the pair-instability mass gap can collapse directly into BH seeds with masses of $\sim 10^{2}$ $\mathrm{M}_{\odot}$ (e.g. Bromm & Larson, 2004; Bromm, 2013; Woosley, 2017; Haemmerlé et al., 2020; Mestichelli et al., 2024). Within the Pop III cluster, stars can collide and form an even more massive star, collapsing into a heavier BH seed (e.g. Katz et al., 2015; Sakurai et al., 2017; Reinoso et al., 2018, 2022; Vergara et al., 2021; Tanikawa et al., 2022a; Wang et al., 2022). Reinoso et al. (2025) shows that IMBHs with masses of up to $10^{4}$ $\mathrm{M}_{\odot}$ are possible to form through stellar mergers and gas accretion. (iii) A pristine gas cloud can avoid fragmentation, collapsing directly into a massive star that eventually forms a heavy seed with masses of around $10^{4}$-$10^{6}$ $\mathrm{M}_{\odot}$ (e.g. Bromm & Loeb, 2003; Begelman et al., 2006, 2008; Begelman, 2010; Chon & Omukai, 2025; van Dokkum et al., 2025). Extremely massive stars with masses of $10^{5}$-$10^{6}\penalty 10000\ \mathrm{M}_{\odot}{}$ can explode by the general relativistic instability supernova mechanism (e.g. Shibata & Shapiro, 2002; Sakurai et al., 2015; Uchida et al., 2017).
@@
As was noted above, metal-poor Population (Pop) III star clusters at high redshifts ($z\gtrsim 10$) with a top-heavy initial mass function (IMF; e.g. Sharda & Krumholz 2022) are prime candidates to produce BH seeds for galactic nuclei and nuclear star clusters (NSCs Neumayer et al., 2020; Greene et al., 2020; Askar et al., 2021, 2022; Schleicher et al., 2022). The IMBHs that form in these clusters can grow even more massive through tidal disruption events over long timescales (Stone et al., 2017; Sakurai et al., 2019; Wang et al., 2025b). The single- and multi-generation mergers from the gravitational runaway process of BHs and other compact objects originating from Pop III star clusters will be relevant gravitational wave (GW) detection events (e.g. Schneider et al., 2000, 2002, 2003; Kinugawa et al., 2014, 2016, 2021a, 2021b; Hartwig et al., 2016; Belczynski et al., 2017; Ng et al., 2022; Mestichelli et al., 2024; Liu et al., 2024a, b; Wu et al., 2025; Wang et al., 2025a). Liu et al. (2025) shows that binary BH mergers in Pop III star clusters can produce GWs, similar to those observed in GW231123. Plunkett et al. (2025) develops observational strategies to constrain the stellar demographics of Pop III stars with next-generation telescopes, especially for the proposed third-generation ground-based GW detectors, the Cosmic Explorer (CE Reitze et al., 2019; Evans et al., 2021), Einstein Telescope (ET Punturo et al., 2010b, a; Sathyaprakash et al., 2012; Maggiore et al., 2020; Liu & Bromm, 2020, 2021), and for future space-borne detectors operating in the millihertz and decihertz frequency bands, such as the laser interferometer space antenna (LISA Amaro-Seoane et al., 2013, 2017, 2023), (TianQin Luo et al., 2016; Liu et al., 2020), or the decihertz GW observatory (DECIGO Kawamura et al., 2011, 2021), where IMBHs are expected to be bright GW sources (e.g. Miller & Hamilton, 2002; Amaro-Seoane et al., 2007; Amaro-Seoane, 2018; Jani et al., 2020; Arca Sedda et al., 2020a, 2021).
@@
Using detailed binary population synthesis (BPS), Tanikawa et al. (2021b) conducted studies on the merger rate density of Pop III binary BHs below, within, and above the pair-instability mass gap from isolated binary stellar evolution. They find that mergers between two low-mass BHs (in their models, low mass means $M<50$ $\mathrm{M}_{\odot}$) independent of mass ratio and semi-major axis distributions of the primordial Pop III binaries could be identified from the BHBH mergers observed by the (Advanced) Laser Interferometer Gravitational-Wave Observatory ((a)LIGO Aasi et al., 2015; Abbott et al., 2018, 2019), (Advanced) Virgo Interferometer ((a)Virgo Acernese et al., 2015; Abbott et al., 2018, 2019) and by extension also the Kamioka Gravitational Wave Detector (KAGRA Abbott et al., 2018, 2020; Kagra Collaboration et al., 2019), although the predicted present-day (10 Gyr) merger rates would be comparatively low ($\sim 0.1\penalty 10000\ \mathrm{yr}^{-1}\mathrm{Gpc}^{-3}$) with regard to the merger rate density inferred by the GWTC-4 catalogue of $15.4^{+6.0}_{-4.5}$ $\mathrm{yr}^{-1}\mathrm{Gpc}^{-3}$ (The LIGO Scientific Collaboration et al., 2025). Similarly, mergers between a low-mass BH and a high-mass BH (in their models, high mass means $M>130$ $\mathrm{M}_{\odot}$ due to the pair instability mass gap) or mergers between two high-mass BHs will be detectable using the aforementioned currently available GW detectors according to Tanikawa et al. (2021b). However, the authors caution that if two conditions hold simultaneously, namely a wide minimum semi-major axis in primordial Pop III binaries due stellar expansion in the protostellar phases and fast stellar rotation that causes excitation of (quasi-)chemically homogeneous evolution of these stars, during which Pop III stars stay more compact than their non-rotating counterparts due to mixed helium throughout the star (e.g. Maeder, 1987; Yoon & Langer, 2005; de Mink et al., 2009). Rapidly rotating Pop III stars (Nandal et al., 2024; Tsiatsiou et al., 2024) have been proposed as a possible source of the high nitrogen abundance observed in galaxies at high redshift (e.g. Bouwens et al., 2010; Tacchella et al., 2023; Marques-Chaves et al., 2024; Naidu et al., 2026). Rotation in Pop III stars induces strong internal mixing, which enlarges the convective core and results in brighter stars, improving the chances of detecting these primordial objects in the early Universe with JWST (Hassan et al., 2025).
@@
Direct detections of Pop III stars, their remnants, or their host clusters remain lacking because they lie at very high redshift. Schauer et al. (2022) reported a candidate extremely massive Pop III star at $z=6.2$, but this requires confirmation. Vanzella et al. (2020) claimed a Pop III complex at $z=6.629$ with the MUSE Deep Lensed Field targeting the Hubble Frontier Field (HFF) galaxy cluster MACS J0416. Using JWST’s NIRCam and lensing, Fujimoto et al. (2025) identified a strong Pop III cluster candidate at $z=6.5$. Another candidate is the high-redshift galaxy RX J2129–z8He, which exhibits He II emission at $1640\mathring$$\mathrm${A}} (Wang et al., 2024). Despite JWST, unambiguous direct observations remain difficult (Rydberg et al., 2013; Katz et al., 2022). de Souza et al. (2013) suggest that several hundred JWST supernova detections could constrain the Pop III IMF. Future wide-field surveys with Euclid (Laureijs et al., 2011; Tanikawa et al., 2022b) and Roman will likely outperform JWST in constraining direct-collapse BHs above the pair-instability gap because they cover larger areas (Lazar & Bromm, 2022; Vikaeus et al., 2022).
@@
Because direct observations of Pop III stars and their environments are lacking, the parameter space for their formation is poorly constrained (e.g. Klessen, 2019; Klessen & Glover, 2023). Fraser et al. (2017) inferred a Pop III IMF from 29 extremely metal-poor Pop II stars that are expected to form in environments enriched by Pop III SNe, and found a Pop I/II-like IMF, but the small sample prevents strong conclusions from being drawn. Theory and simulations generally predict top-heavy Pop III IMFs; such IMFs favour XRB populations with brighter and more spatially irregular X-ray emission (Sartorio et al., 2023). Hydrodynamic models are therefore widely used to probe plausible Pop III IMFs and their environmental dependence (e.g. Hirano & Bromm, 2017, 2018; Susa, 2019; Chon et al., 2021; Sharda et al., 2021; Latif et al., 2022). The binary fraction and initial binary statistics remain highly uncertain (Stacy & Bromm, 2013; Liu et al., 2021a) and, being environment-dependent, affect massive-star and compact-object demographics (Hirano et al., 2018; Sugimura et al., 2020; Kroupa, 2025).
@@
The observational and theoretical uncertainties are inherited by the initial conditions of Pop III star cluster simulations. One such parameter concerns the degree of initial Pop III star cluster rotation. A set of direct N-body simulations of initially rotating Pop III star clusters were performed by Vergara et al. (2021, $N=10^{3}-10^{4}$ and evolution only up to $2\penalty 10000\ \mathrm{Myr}$). They used Miyamoto-Nagai models with flattening and rotation and found that not only the collision rate increases with increasing bulk rotation, but additionally the number of escapers is reduced the larger the initial rotation is. The low particle numbers, the short simulation time, and other simplifying assumptions, such as neglecting stellar evolution, make it difficult to generalize these results. In general, for simulations in collisional dynamics of rotating star clusters, distribution functions from, for example, Goodman (1983), Longaretti & Lagoute (1996), Einsel & Spurzem (1999), and Varri & Bertin (2012) are typically used. Kamlah et al. (2022b) recently ran simulations of Pop II star clusters with rotating King models from Einsel & Spurzem (1999, using the 2D Fokker-Planck code Fopax; ). Apart from the formation and dissolution of a rotating bar of BHs that is related to the gravothermal-gravogyro catastrophe (Hachisu, 1979, 1982; Akiyama & Sugimoto, 1989; Einsel & Spurzem, 1999; Hong et al., 2013; Kamlah et al., 2022b), they found a possible dependence of BHBH binary abundances on the initial star cluster bulk rotation; for a fast-rotating model (rotating King model $W_{0}=6.0,\penalty 10000\ \omega_{0}=1.2$), they found significantly more BHBH binaries during a phase of star cluster evolution when BHBH mergers in the aforementioned ‘slow’ regime are relevant (Greene et al., 2020). Similarly, Webb et al. (2019) found that increased initial star cluster rotation precipitates the formation of circularized BHBH binaries. Since Pop III are postulated to harbour many more (IM)BHs than Pop II star clusters, relative to their size, we tested whether earlier findings apply to Pop III clusters.
@@
Furthermore, the lack of concrete observations of Pop III stars adds uncertainty to their stellar evolution. The most impactful parameter that differentiates stars of the same masses in Pop I and Pop III populations is the metallicity (therefore, Pop III stars are also referred to as extreme metal poor stars e.g. in Tanikawa et al., 2020). The lower the metallicity, the weaker the radiation-driven winds that affect massive stars. Generally, this statement also holds for pulsation-driven winds (Nakauchi et al., 2020). In the regime of $Z/\mathrm{Z}_{\odot}{}<10^{-4}$ both winds become negligibly small. However, it has been suggested that Pop III stars form with very high rotation rates, when the magnetic fields are negligible, and thus there is negligible magnetic braking (Stacy et al., 2011, 2013; Hirano & Bromm, 2018). The winds due to the weakened stellar magnetic field of such stars may be very powerful (Liu et al., 2021b; Jeena et al., 2023).
@@
The lower the metal in a Pop III star, the more drastic its internal evolutionary changes, and the lower its opacity (Ekström et al., 2008). First, the lower the metallicity of a star, the more compact it is, because the line-driven radiation pressure diminishes. Furthermore, most massive Pop I/II stars are characterized by a red supergiant (RSG) evolutionary phase with convective envelopes, while most Pop III stars end with a blue supergiant (BSG) phase and radiative envelopes (Tanikawa et al., 2020). These properties affect binary stellar evolution. While Pop I/Pop II stars tend to undergo unstable mass transfer and common envelope evolution (CEE) in the RSG phase, BSGs from Pop III stars undergo stable mass transfer, so that less mass is ejected from the binary system. As a result, BHBH binaries from Pop III stellar populations can be more massive than Pop I/II counterparts, even ignoring wind mass loss, making Pop III star clusters a very attractive target of GW event detection and progenitor studies (e.g. Inayoshi et al., 2017; Kinugawa et al., 2021a, b; Tanikawa et al., 2021b, a, 2022c, 2022a; Liu et al., 2024a). Dense stellar environments substantially increase the fraction of mergers involving BHs with primary masses exceeding the pair-instability mass limit (Mestichelli et al., 2024, 2025). Furthermore, shorter initial orbital periods lead to higher merger rates Santoliquido et al. (2023), which can boost the production of GW sources in the early Universe.
@@
In this paper, we present eight direct N-body simulations of extremely massive, rotating Pop III clusters, run with and without self-consistent general relativistic (GR) merger recoil kicks. The main novelty is the inclusion of initial bulk rotation. We examine how initial bulk rotation, state-of-the-art Pop III stellar-evolution models, GR merger recoil kicks (GRKs), primordial binaries, a continuous IMF, and tidal-field mass loss jointly shape cluster dynamics and IMBH formation.
@@
The paper is structured as follows. Section 2 summarizes the methodology, the implementation of the GRKs, and the Pop III stellar evolution fitting formulae. Section 3 outlines the initial conditions. Section 4 presents the results. In Sect. 5 we summarize and discuss the work.